OpenStudy (anonymous):

Can you find the square root of √-16

OpenStudy (kropot72):

The imaginary number i = square root of -1. Extracting -1 from under the radical sign gives: \[\sqrt{-16}=i \sqrt{16}=?\]

OpenStudy (anonymous):

But can you find the square root of it?

OpenStudy (anonymous):

The square root of 16 s 4 because 4*4=16

OpenStudy (kropot72):

@FastLearner1234 The square root of the expression in your question can be found only by expressing the solution as an imaginary number.

OpenStudy (anonymous):

thanks you are very helpful what grade are you in..smarty person lol

OpenStudy (kropot72):

You're welcome. I'm not in grades anymore :)